# Generalized Simple Graphical Rules for Assessing Selection Bias

### v1

@@ Abstract
Keywords: selection bias, causal inference, causal diagrams, SWIG, inverse probability weighting
@@ 1 Introduction
Selection bias is one of the fundamental obstacles to causal inference. The concept of selection bias encompasses various forms stemming from different mechanisms for individuals being selected into the study or analysis [1]. The complexity of these selection mechanisms poses difficulty in a structural definition of selection bias, which has been continually evolving over the past two decades [2, 3, 4, 5]. Recently, a refined definition of selection bias that unifies different types of selection bias was proposed [3] by considering the true causal effect in the population prior to the selection process as the reference. This definition can be decomposed into two types: type 1 selection bias due to restricting to at least one level of a collider or its descendant, and type 2 selection bias due to restricting to at least one level of an effect measure modifier.
@@
With a clearer understanding of selection bias in place, recent methodological developments have also been dedicated to recovering causal effects in the presence of selection bias. Specifically, based on the fundamentals of causal diagrams, including directed acyclic graphs (DAGs) and single-world intervention graphs (SWIGs), simple graphical rules were proposed to account for selection bias, alone or possibly combined with other issues (e.g., confounding bias or missing data) [6, 7, 8, 9]. These simple graphical rules are usually coupled with specific simple identification strategies and estimators, allowing researchers to recover causal effects in some identifiable cases without the need to search for alternative formulas.
@@
However, there are still important cases of selection bias unaddressed by common graphical rules proposed previously. For example, in the graphical rule proposed by Mathur and Shpitser, one premise is that post-treatment variables can not be conditioned on in their selected-sample analysis estimator [8]. Additionally, they focus on the conditional average treatment effect (CATE) rather than the average treatment effect (ATE). This limits the use of external information to recover the ATE if it is of interest. Some other simple rules do not prohibit adjustments for post-treatment variables, but they are rather stringent to avoid confounding bias [10, 7, 6].
@@
In some scenarios, the covariate information available to address selection bias consists of post-treatment variables [11]. In these cases, the identification of ATE is still possible despite the rejection by these specific rules. This highlights the importance of applying fundamental graphical rules to search for identification strategies. For example, approaches such as inverse probability weighting (IPW) methods and g-computation are powerful tools for recovering treatment effects if external information beyond the selected sample is available [1, 7, 11].
@@
In this article, we discuss two important cases of selection bias unaddressed by existing simple graphical rules. The first one represents the selection bias where a descendant of a collider is being conditioned on. The other is an example where sample selection is affected by the mediator. We prove that the average treatment effect (ATE) in both cases can be recovered by the g-computation or IPW methods if the external information on the covariates is available in the entire population. The conditions for proving these identification strategies can also be considered graphical rules for addressing selection bias in similar cases. We propose two specific IPW estimators and their variance estimators for the ATE of interest. The necessary steps are outlined to explain the practical procedures. We conduct simulation studies to verify the statistical properties of the IPW estimators. In our hypothetical but realistic simulation settings, we show that the conventional crude complete-case selected-sample analysis yields conclusions that are contradictory to the truth, potentially misleading decision-making.
@@ 2 Two cases of selection bias unaddressed by simple graphical rules
It is difficult to address selection bias because of numerous combinations of graphical structures and identification formulas. As a result, researchers usually consider some simple graphical rules and a particular class of identification formulas and define ‘identifiability’ (or ‘recoverability’) in a narrow sense of whether the treatment effect of interest can be identified by that class of formulas [10, 7, 6, 8]. Simple graphical rules are usually sufficient conditions for verifying that the treatment effect can be identified (or recovered) by a particular class of identification formulas. If these conditions are violated, it does not rule out the possibility that the effect of interest can still be identified by other formulas. In this section, we show two important cases of selection bias unaddressed by the previously proposed simple graphical rules. The identification of treatment effects for similar cases by alternative formulas will be discussed in later sections.
@@ 2.1 Setting
Both Case 1 and Case 2 are important as they represent scenarios that occur in real-world settings. For example, consider Case 1, where the study is a randomized trial of a Prophylactic medication for malaria among susceptibles. The treatment $A$ is a binary variable indicating assignment to the medication or not. The outcome of interest $Y$ is the incidence of malaria or not. The incidence of fever or not, $L$, directly affects the receipt of the blood test for malaria or not, $S$, for the outcome to be detected. It was suggested that the selection bias in this causal graph can not be addressed to recover the ATE by the selection-backdoor criteria [6].
In Case 2, the study is a randomized trial of a vaccine for an infectious disease with multiple endpoints. The treatment $A$ indicates assignment to the vaccine or not. The outcome $Y$ is the symptom of having a headache or not. The outcome is mediated by the incidence of infection or not, $L$, which also affects whether a unit reports the outcome status or not, $S$.
@@
Figure 1: Causal directed acyclic graphs (DAGs) and single-world intervention graphs (SWIGs) for Case 1 and Case 2 of sample selection directly affected by a post-treatment variable $L$. When a unit is selected into the sample (i.e. $S_{i}=1$), we observe all variables $(A_{i},Y_{i},L_{i})$. When a unit is not selected (i.e. $S_{i}=0$), external information on $L_{i}$ among unselected units, along with the knowledge of $A_{i}$ or the rule of treatment assignment $p(A_{i}=1)$, might still be available. In Case 1, the direct upstream cause of selection, $L$, is the common effect (i.e., collider) of both the treatment $A$ and the outcome $Y$. In Case 2, the direct upstream cause of selection, $L$, is the mediator of the treatment $A$ on the outcome $Y$.
@@ 2.2 The failure by existing simple graphical rules
We focus on three existing simple graphical rules (i.e., the selection-backdoor criterion [6], the generalized adjustment criterion [7], and Mathur and Shpitser’s graphical rules [8]) proposed over the past decade. In Appendix A, we show that the two cases as discussed above violate these existing simple graphical rules, and the identification by the corresponding formulas is not guaranteed.
@@ 3 Identification by the fundamentals of graphical rules
Existing simple graphical rules are derived as compared to and based on the fundamentals of DAGs and SWIGs [12, 13]. The core element is to read off conditional independence based on the fundamental definition of d-separation that invites an open-ended search for an identification formula. When the graphical structure of the problem falls outside of the expertise by the existing simple graphical rules, it naturally invites an ad-hoc investigation of the identifiability and possible construction of unconventional formulas. If the ATE can be proved to be identified by some other formulas the conditions that validate the identification can be extracted as new graphical rules for these formulas. Their utility depends on how often the stipulated conditions hold under realistic circumstances.
In this section, we first introduce a g-computation formula that was acknowledged to be successful in identifying the ATE by conditioning on post-treatment variables [11]. Then we propose another inverse probability weighting (IPW) method to identify the ATE that addresses selection bias in the two cases discussed. The identification results by the two proposed methods is generalized under specific graphical rules to allow (i) observed confounders for the relationships between the treatment and post-treatment variables, (ii) unobserved confounders for the relationships between post-treatment variables and the outcome, and (iii) selection to be directly affected by the treatment.
@@ 3.1 G-computation formulas
A special-case g-formula conditioning on a post-treatment variable $Z$ was derived for identifying the ATE in a practical example of randomized controlled trials, demonstrating the utility of SWIGs [11]. We extend their results by allowing the existence of confounders in an observational setting.
We propose the following theorem to conclude the general results for its application.
@@
Therefore, despite being motivated by an ad-hoc estimator for a special case, the g-formula (1) recovers the ATE when some post-treatment variables can be conditioned on to address selection bias. The conditions can be viewed as derived graphical rules in this specific context. The example by Breskin et al. [11] is a special case of our theorem (where they denote the post-treatment variable by $W$ instead of $L$) by the following corollary.
@@
Taking $Z=L$, our Case 2 was explicitly mentioned as an extension to the graph in Breskin et al. [11] where the derivation of the formula applies. Moreover, it is straightforward to verify that the conditions (2) are satisfied in both of our cases by the SWIGs in Fig. 1. We have the following corollary as examples.
@@ 3.2 Inverse probability weighting
The first condition assumes that all paths between $Y(a)$ and $S(a)$ are blocked by $Z(a)$ and $X$. The next two conditions assume that the treatment and the potential responses of post-treatment variables (including the outcome, covariates, and selection) can be sequentially separated. The second condition for the IPW method (4) is slightly different from that for the g-formula (2). The requirement by the IPW method that $S(a)$ should be conditionally separated from $A$ is due to the inclusion of the inverse probability weighting model for the selection to identify the ATE.
@@
The treatment-selection confounders $X_{3}$ cannot be omitted by the IPW method for identifying the ATE because the model for the selection deems it necessary to collect these variables for correct causal model specification. This is different from the requirements by the g-formula for the same graph.
@@ 5 Simulation
We conduct simulation studies to verify the performance of our IPW estimators in both cases of Fig. 1. The null hypothesis of the treatment effect is set to be true in the first case, but not in the second case. For each case, we generate datasets with varying sample sizes of $n=100,n=1000,n=10000$. The two IPW estimators are compared to the crude selected-sample analysis which only includes the observed treatment $A$ in the simple linear regression of the observed outcome $Y$. In each setting, we simulate $K=10000$ datasets and fit the models. Performance measures including average point estimates, mean squared error, confidence intervals, coverage rates of the true effect, and rejection rates of the null hypothesis, are reported as results. We show that the crude analysis yields biased and misleading conclusions but our IPW methods accurately recover the true effect and enable the correct decision from its inference.
@@
Figure 3: Case 1: simulation results of average point estimates of ATE, mean squared error, confidence intervals, coverage rates of the true ATE by the 95% confidence interval, and the type I error rate by the crude selected-sample complete-case analysis and the two proposed IPW estimators under different sample sizes. The true ATE is 0 for Case 1.
The crude selected-sample analysis (i.e., crude complete-case analysis) in Case 1 (red colored in Fig. 3) erroneously concludes a beneficial effect of the medication on reducing the risk of malaria. The $95\%$ confidence interval only contains the truth when the sample size is small due to huge uncertainty. The coverage rates of the truth, zero effect, by its $95\%$ confidence interval drops to zero as the sample size increases. In this null effect case, it implies that the rejection rate rises to $100\%$ as the sample size increases. The intuition is that, with a low probability of measuring the outcome among those who do not have a fever, individuals who are not assigned to the medication and do not have malaria are unlikely to be selected. Without information from these healthy individuals, the effect of the medication is overestimated, leading to selection bias. As a result, the crude selected-sample analysis yields a misleading result by exaggerating the impact of the medication.
The results by the IPW estimators are summarized for Case 1 in Fig. 3 (green and blue colored). Point estimates by both methods are correctly centered around zero. The mean squared error quickly drops and approaches zero. The coverage rates of the truth by the $95\%$ confidence intervals converge towards $95\%$ as the sample size increases. Equivalently, the type I error rate of rejecting the null hypothesis approaches $5\%$. Both methods are expected to properly fail to reject the null hypothesis at the intended size. The Hájek estimator has slightly smaller variances and narrower confidence intervals than the HT estimator.
In Case 2, consider a randomized trial of some vaccine for an infectious disease that assigns half of the individuals to the treatment group. One of the endpoints in this trial is the risk of having a headache within a certain period. The treatment ($A$) is random assignment to the vaccine or the placebo group. Its effect on the outcome of having a headache or not ($Y$), can be directly through its side effects by increasing the risk. It can also be indirectly, mediated by the risk of being infected. The incidence of infection or not ($L$) affects the chance that they report whether they had a headache or not ($S$) during that period. Infected individuals are more likely to have a hospital encounter and report their symptoms in terms of headache, but those uninfected are less likely to report having headache symptoms or not unless through follow-up contacts. The data-generating process is
@@
Figure 4: Case 2: simulation results of average point estimates of ATE, mean squared error, confidence intervals, coverage rates of the true ATE by the 95% confidence interval, and the type II error rate by the crude selected-sample complete-case analysis and the two proposed IPW estimators under different sample sizes. The true ATE is -0.15 for Case 2.
The crude selected-sample complete-case analysis of Case 2 (red colored in Fig. 4) exhibits biased estimates of the ATE as small as $-0.03$. The $95\%$ confidence intervals contain zero across all three settings. Similar to Case 1, the coverage rate of the true ATE also drops to zero when the sample size increases. The rejection rate is lower than $10\%$ when the sample size is $1000$ or smaller, indicating a type II error rate higher than $90\%$. When the sample size reaches $1000$, the type II error rate is still higher than $50\%$. The explanation is, that the decreased chance of reporting whether an individual has a headache or not among those uninfected, partly masks the beneficial indirect effect of the vaccine on reducing headaches through reducing their risk of infection. Therefore, the crude selected-sample analysis fails to reveal the actual beneficial impact of the vaccination on reducing the overall risk of headaches.
Tabel 4 shows the results of the IPW estimators for Case 2 (colored in green and blue). Both methods successfully recover the true ATE around $-0.15$ across all sample sizes. The mean squared error converges to zero. The coverage rates of the true ATE by the $95\%$ confidence interval increase to the intended level. The rejection rates increases to $100\%$ when the sample size is large enough. The power of the two methods with a sample size of $1000$ is greater than that of the crude complete-case analysis with a sample size of $10000$ by over $10\%$.

### v3

@@ Abstract
Keywords: selection bias, causal inference, causal diagrams, SWIG, inverse probability weighting, epidemiologic research
@@ 1 Introduction
Selection bias is one of the fundamental obstacles to causal inference. The concept of selection bias encompasses various forms stemming from different mechanisms for individuals being selected into the study or analysis [1]. The complexity of these selection mechanisms poses difficulty in a structural definition of selection bias, which has been continually evolving over the past two decades [2, 3, 4, 5]. Recently, a refined definition of selection bias that unifies different types of selection bias was proposed [3] by considering the true causal effect in the population prior to the selection process as the reference. This definition can be decomposed into two types: type 1 selection bias due to restricting to at least one level of a collider or its descendant, and type 2 selection bias due to restricting to at least one level of an effect measure modifier.[3]
@@
With a clearer understanding of selection bias in place, recent methodological developments have also been dedicated to recovering causal effects in the presence of selection bias. Specifically, based on the fundamentals of causal diagrams, including directed acyclic graphs (DAGs) and single-world intervention graphs (SWIGs), simple graphical rules were proposed to account for selection bias, alone or possibly combined with other issues (e.g., confounding bias or missing data) [6, 7, 8, 9].
@@
However, there are still important cases of selection bias that cannot be addressed by these existing graphical rules. For example, in the graphical rules proposed by Mathur and Shpitser, one premise is that post-treatment variables cannot be conditioned on in their selected-sample analysis estimator [8]. In addition, they focus on the conditional average treatment effect (CATE) rather than the average treatment effect (ATE, which is arguably more widely used). Some other existing graphical rules do not prohibit adjustments for post-treatment variables, but they are rather stringent to avoid confounding bias [10, 7, 6].
@@
In some cases, the covariate information available to address selection bias may only include post-treatment variables [11]. In these cases, the identification of ATE may still be possible despite the rejection by the existing graphical rules. This highlights the importance of applying fundamental graphical rules to search for identification strategies. For example, approaches such as inverse probability weighting (IPW) methods and g-computation are powerful tools for recovering treatment effects if external information beyond the selected sample is available [1, 7, 11].
@@
In this article, we discuss two important cases of selection bias unaddressed by existing graphical rules. The first one represents the selection bias where a descendant of a collider of the exposure and the outcome is being conditioned on. The other is an example where sample selection is affected by the mediator. We prove that the average treatment effect (ATE) in both cases can be recovered by the g-computation or IPW methods if the external information on the covariates is available in the general population. The conditions for proving these identification strategies are provided and can also be used for addressing selection bias in similar cases. We propose two specific IPW estimators and their variance estimators for the ATE. We conduct simulation studies to verify the statistical properties of the proposed IPW estimators. In our hypothetical simulation settings, we show that the conventional crude complete-case analysis using the selected sample yields conclusions contradictory to the truth and can mislead decision-making.
@@ 2 Two cases of selection bias unaddressed by simple graphical rules
To detect and control for selection bias, researchers usually consider some simple graphical rules along with a specific class of identification formulas, and then define ‘identifiability’ (or ‘recoverability’) in a narrow sense of whether the treatment effect of interest can be identified by that class of formulas [10, 7, 6, 8]. Simple graphical rules are usually sufficient conditions for verifying that the treatment effect can be identified (or recovered) by a particular class of identification formulas. If these conditions are violated, it does not rule out the possibility that the effect of interest can still be identified by other formulas. In this section, we show two important cases of selection bias that cannot be addressed by the previously proposed graphical rules. The identification of treatment effects for similar causal structures by modified formulas will be discussed in later sections.
@@ 2.1 Setting
Both Case 1 and Case 2 are important as they represent scenarios that may occur in real-world settings. For example, consider Case 1, where the study is a randomized trial of a prophylactic medication for malaria among susceptibles. The treatment $A$ is a binary variable indicating assignment to the medication or not. The outcome of interest $Y$ is the incidence of malaria or not. The incidence of fever or not, $L$, directly affects the receipt of the blood test for malaria or not, $S$, for the outcome to be detected. It was suggested that the selection bias in this exact causal graph cannot be addressed to recover the ATE by the selection-backdoor criteria [6].
In Case 2, the study is a randomized trial of a vaccine for an infectious disease with multiple endpoints. The treatment $A$ indicates assignment to the vaccine or not. The outcome $Y$ is the symptom of having a headache or not. The outcome is mediated by the incidence of infection, $L$, which also affects whether a unit reports the outcome status, $S$.
@@
Figure 1: Causal directed acyclic graphs (DAGs) and single-world intervention graphs (SWIGs) for Case 1 and Case 2 of sample selection directly affected by a post-treatment variable $L$. When a unit is selected into the sample (i.e. $S_{i}=1$), we observe all variables $(A_{i},Y_{i},L_{i})$. When a unit is not selected (i.e. $S_{i}=0$), external information on $L_{i}$ among unselected units, along with the knowledge of $A_{i}$ or the rule of treatment assignment $p(A_{i}=1)$, might still be available, but the outcome $Y_{i}$ is unobserved. In Case 1, the direct upstream cause of selection, $L$, is the common effect (i.e., collider) of both the treatment $A$ and the outcome $Y$. In Case 2, the direct upstream cause of selection, $L$, is the mediator of the treatment $A$ on the outcome $Y$.
@@ 2.2 The failure by existing simple graphical rules
We focus on three existing simple graphical rules (i.e., the selection-backdoor criterion [6], the generalized adjustment criterion [7], and Mathur and Shpitser’s graphical rules [8]) proposed over the past decade. In Appendix A, we show that the two cases as discussed above violate these existing simple graphical rules, and the identification by their corresponding formulas is not guaranteed.
@@ 3 Identification by the fundamentals of graphical rules
In this section, we first introduce a g-computation formula that can be used to recover the ATE by conditioning on post-treatment variables [11]. Then we propose the inverse probability weighting (IPW) method to recover the ATE and address selection bias in the two cases discussed. The identification results by the two proposed methods are generalized under specific graphical rules to allow (i) observed confounders for the relationships between the treatment and post-treatment variables, (ii) unobserved confounders for the relationships between post-treatment variables and the outcome, and (iii) selection that is directly affected by the treatment.
@@ 3.1 G-computation formulas
In Breskin et al.[11], a special-case g-formula conditioning on a post-treatment variable $Z$ was derived for identifying the ATE in a practical example of randomized controlled trials, demonstrating the utility of SWIGs. We extend their results by allowing the existence of confounders in an observational setting.
Specifically, we propose the following theorem.
@@
Therefore, despite being motivated by an ad-hoc estimator for a special case, the g-formula (1) will recover the ATE when some post-treatment variables can be conditioned on to address selection bias. The conditions can be viewed as derived graphical rules in this specific context. The example by Breskin et al. [11] is a special case of our theorem (where they denote the post-treatment variable by $W$ instead of $L$) by the following corollary.
@@
Taking $Z=L$, our Case 2 can be considered as a modification of the graph in Breskin et al. [11] where the derivation of the formula still applies. Moreover, it is straightforward to verify that the conditions (2) are satisfied in both of our cases by the SWIGs in Fig. 1. We have the following corollary as examples.
@@ 3.2 Inverse probability weighting
The first condition assumes that all paths between $Y(a)$ and $S(a)$ are blocked by $Z(a)$ and $X$. The next two conditions assume that the treatment and the potential responses of post-treatment variables (including the outcome, covariates, and selection) can be sequentially separated. The second condition for the IPW method (4) is slightly different from that for the g-formula (2). The requirement by the IPW method that $S(a)$ should be conditionally separated from $A$ is due to the inclusion of the inverse probability weighting model for the selection to recover the ATE.
@@
The treatment-selection confounders $X_{3}$ cannot be omitted by the IPW method for recovering the ATE because the model for the selection deems it necessary to collect and adjust for these variables for correct causal model specification. This is different from the requirements by the g-formula for the same graph.
@@ 5 Simulation
We conduct simulation studies to verify the performance of our IPW estimators in both cases of Fig. 1. The null hypothesis of the treatment effect is set to be true in the first case, but not in the second case. For each case, we generate datasets with varying sample sizes of $n=100,n=1000,n=10000$. The two IPW estimators are compared to the crude selected-sample analysis which only includes the observed treatment $A$ in the simple linear regression of the observed outcome $Y$. In each setting, we simulate $K=10000$ datasets and fit the models. We show that the crude analysis yields biased and misleading conclusions but our IPW methods accurately recover the true effect.
@@
Figure 3: Case 1: simulation results of average point estimates of ATE, mean squared error, confidence intervals, coverage rates of the true ATE by the 95% confidence interval, and the type I error rate by the crude selected-sample analysis using the selected sample and the two proposed IPW estimators under different sample sizes. The true ATE is 0 for Case 1.
The crude selected-sample analysis (i.e., crude complete-case analysis) in Case 1 (red colored in Fig. 3) erroneously concludes a beneficial effect of the medication on reducing the risk of malaria. The $95\%$ confidence interval only contains the truth when the sample size is small due to substantial uncertainty. The coverage rates of the truth, zero effect, by its $95\%$ confidence interval drops to zero as the sample size increases. In this null effect case, it implies that the rejection rate rises to $100\%$ as the sample size increases. The intuition is that, with a low probability of measuring the outcome among those who do not have a fever, individuals who are not assigned to the medication and do not have malaria are unlikely to be selected. Without information from these healthy individuals, the effect of the medication is overestimated, leading to selection bias. As a result, the crude selected-sample analysis yields a misleading result by exaggerating the impact of the medication.
The results by the IPW estimators are summarized for Case 1 in Fig. 3 (green and blue colored). Point estimates by both methods are correctly centered around zero. The mean squared error quickly drops and approaches zero. The coverage rates of the truth by the $95\%$ confidence intervals converge towards $95\%$ as the sample size increases. Equivalently, the type I error rate of rejecting the null hypothesis approaches $5\%$. Both methods are expected to fail to reject the null hypothesis at the intended size. The Hájek estimator has slightly smaller variances and narrower confidence intervals than the HT estimator.
In Case 2, consider a randomized trial of some vaccine for an infectious disease that assigns half of the individuals to the treatment group. One of the endpoints in this trial is the risk of having a headache within a certain period. The treatment ($A$) is random assignment to the vaccine or the placebo group. Its effect on the outcome of having a headache or not ($Y$), can be directly through its side effects by increasing the risk. It can also be indirectly, mediated by the risk of being infected. The incidence of infection ($L$) affects the chance that they report whether they had a headache ($S$) during that period. Infected individuals are more likely to have a hospital encounter and report their symptoms in terms of headache, but those uninfected are less likely to report having headache symptoms or not unless through follow-up contacts. The data-generating process is
@@
Figure 4: Case 2: simulation results of average point estimates of ATE, mean squared error, confidence intervals, coverage rates of the true ATE by the 95% confidence interval, and the type II error rate by the crude complete-case analysis using the selected sample and the two proposed IPW estimators under different sample sizes. The true ATE is -0.15 for Case 2.
The crude selected-sample analysis of Case 2 (red colored in Fig. 4) exhibits biased estimates of the ATE as small as $-0.03$. The $95\%$ confidence intervals contain zero across all three settings. Similar to Case 1, the coverage rate of the true ATE also drops to zero when the sample size increases. The rejection rate is lower than $10\%$ when the sample size is $1000$ or smaller, indicating a type II error rate higher than $90\%$. When the sample size reaches $1000$, the type II error rate is still higher than $50\%$. The explanation is that the decreased chance of reporting whether an individual has a headache among those uninfected partly masks the beneficial indirect effect of the vaccine on reducing headaches through reducing their risk of infection. Therefore, the crude selected-sample analysis fails to reveal the actual beneficial impact of the vaccination on reducing the overall risk of headaches.
Tabel 4 shows the results of the IPW estimators for Case 2 (colored in green and blue). Both methods successfully recover the true ATE around $-0.15$ across all sample sizes. The mean squared error converges to zero. The coverage rates of the true ATE by the $95\%$ confidence interval increase to the intended level. The rejection rates increases to $100\%$ when the sample size is large enough. The power of the two methods with a sample size of $1000$ is greater than that of the crude selected-sample analysis with a sample size of $10000$ by over $10\%$.